# Hier-SLAM++: Neuro-Symbolic Semantic SLAM with a Hierarchically Categorical Gaussian Splatting

### v1

@@ III-C Monocular setting with geometric priority
Accurate depth information provides valuable geometric cues for global map reconstruction and pose estimation in a SLAM system. In an RGB-D setting, depth signals are directly obtained from sensors. However, in a monocular setting without depth information, SLAM [28, 29, 30, 31] typically relies on triangulation to initialize the global mapping. Yet, due to the inherent coupling between pose estimation and mapping, this approach often suffers from relatively lower accuracy compared with RGB-D based methods. In our Hier-SLAM++, we employ a feed-forward method as a geometric prior. Specifically, we leverage DUSt3R [27] to provide depth geometric priors for monocular SLAM. DUSt3R is able to generate both a 3D point map and pose estimations from several sparse-view image inputs, which can serve as the geometric prior for the monocular SLAM. The following section provides details on the monocular setting in our Hier-SLAM++.
@@ III-C1 Geometric prior acquisition
Each image set is then fed into DUSt3R to generate geometric information for every image frame, specifically the depth maps. Although DUSt3R is also able to provide pose predictions for each input frame, we observe that the accuracy of estimated pose is relatively low—performing even worse than a constant velocity pose prediction in the SLAM system based on our experiments. Therefore, only the estimated geometric information from DUSt3R is leveraged in our method.
@@ III-C2 Geometric prior correction
The scale of the estimation for each image set varies and also differs from the scale of the online global map estimation during SLAM operation, requiring an additional geometric prior correction operator.
Firstly, we calculate the initial scale and shift parameters by comparing the rendering results from the current global map prediction with the depth prior obtained from DUSt3R. Specifically, the depth map $D_{c}$ and the silhouette map $S_{c}$ for current frame $I_{c}$ are rendered based on the global Gaussian map $G$ as well as the pose estimation $\mathbf{T}_{c}$ for the current frame:
@@
where $\mathcal{R}$ represents the rasterization procedure for the 3D Gaussian map. Next, the global scale and shift parameters $\lambda_{c},\tau_{c}$ are calculated by comparing the predicted depth $\hat{D}_{c}$ from DUSt3R with the rendered depth $D_{c}$ using the least squares estimation on the visible regions $M_{c}$ defined by $S_{c}$:
@@
where $\hat{D}_{c}^{\text{align}}=\lambda_{c}\cdot\hat{D}_{c}+\tau_{c}$.
Secondly, the estimated parameters are fed into the optimization as initialization and further optimized jointly within the map optimization via the depth supervision loss:
$\{\lambda_{c},\tau_{c},G\}=\arg\min|\hat{D}_{c}^{\text{align}}-D_{c}|M_{c}$ (17)
In this way, the depth prediction can be continuously refined online with increasing observations, providing geometric supervisory signals during SLAM optimization. This helps guide global reconstruction and pose estimation, leading to improved results throughout the SLAM process.
@@ III-D Semantic Gaussian Splatting SLAM
The global semantic 3D Gaussian Splatting SLAM system is illustrated in the left part of Fig. 2. Based on the established semantic 3D Gaussian representation, global 3D reconstruction (Mapping) and pose estimation (Tracking) are alternately performed with the video stream input. For the RGB-D setting, depth information for each image frame is obtained directly from depth sensors. For the monocular setting, we employ a feed-forward method to obtain the geometric prior, as introduced in the previous section. Further details of our Hier-SLAM++ system are provided in the following section.
@@ III-D1 Semantic 3D Gaussian representation
The semantic 3D Gaussian representation is represented as the Gaussian primitives with the hierarchical semantic embedding. Motivated by [16], we adopt the isotropic and view-independent Gaussian integrated with the semantic embeddings. Each semantic Gaussian is composed of color $\bm{c}$, the center position $\bm{\mu}$, the radius $r$, the opacity $o$, and its semantic embedding $\bm{h}$. And the influence of each Gaussian $G(\bm{X})$ [13] according to the standard Gaussian equation is:
@@
Following [13], every semantic 3D Gaussian primitive within the current global map is projected to the 2D image space using the tile-based differentiable $\alpha$-compositing rendering. The semantic map $H$ is rendered as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{H=\sum_{i=1}^{n}\bm{h}_{i}\alpha_{i}(\bm{X})T_{i}\quad\text{with}\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{X}))}}$ (19)
@@
Different from previous work [16], which uses separate forward and backward Gaussian modules for different parameters, our Hier-SLAM++ system adopt a unified forward and backward modules that processes all parameters, including semantics, color, depth, and silhouette images—significantly improving overall runtime efficiency.
@@ III-D2 Tracking
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbf{T}_{c}=\arg\min_{\mathbf{T}_{c}}L_{\text{Track}}(G,\mathbf{T}_{c})}}$ (21)
@@ III-D3 Mapping
The global map information, including semantic information, is optimized in the mapping procedure while keeping camera poses fixed:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{G=\arg\min_{G}L_{\text{Map}}(G,\mathbf{T}_{i})}}$ (23)
@@ IV-A Experiment settings
The experiments are conducted on both synthetic and real-world datasets, including ScanNet [40], Replica [41], and TUM-RGBD [42]. Following the evaluation metrics which used in previous SLAM methods [16, 33], we leverage ATE RMSE (cm) to quantify SLAM tracking accuracy. For mapping evaluation, Depth L1 (cm) is used to measure accuracy. To assess image rendering quality, we adopt PSNR (dB), SSIM, and LPIPS as evaluation metrics. In line with previous works [24, 7, 11], we assess semantic rendering performance using mIoU (mean Intersection over Union across all classes) to represent global semantic information. To evaluate computational efficiency, we report the runtime of the proposed SLAM method. For comparison, we benchmark our approach against state-of-the-art dense visual SLAM methods, including both NeRF-based and 3D Gaussian SLAM approaches, to demonstrate its effectiveness. Additionally, we compare with state-of-the-art semantic SLAM methods, covering both NeRF-based and Gaussian-based techniques, to highlight its capability in hierarchical semantic representation and scalability. The experiments are conducted in the Nvidia L40S GPU.
For experimental settings, the semantic embedding of each Gaussian primitive is initialized randomly. The semantic optimization loss weights $\omega_{1}$ and $\omega_{2}$ are set to $1.0$ and $0.0$, respectively, for the first $\eta$ iterations, where $\eta$ is set as $15$. Afterwards, $\omega_{1}$ remains $1.0$, while $\omega_{2}$ is increased to $5.0$. This strategy means that we first apply the Inter-level loss to initialize the hierarchical semantic coding, followed by incorporating the Cross-level loss to refine the whole embedding. For tracking loss, we use $\delta=0.99$, with weights $w_{1}=1.0$ and $w_{2}=0.5$. For mapping, the loss weights are set as follows: $w_{3}=1.0$, $w_{4}=0.5$, and $w_{5}=0.2$.
@@
Best results are highlighted as FIRST, SECOND.
@@ IV-B Camera Tracking Accuracy
Tab. I presents our tracking performance on the Replica dataset [41] in the RGB-D setting, compared with state-of-the-art dense SLAM methods, both with and without semantic information. By accurately learning semantic information, our Hier-SLAM++ outperforms all existing methods in 6 out of 8 sequences. For the remaining two sequences, our method achieves nearly the same performance with sub-millimeter level localization error. The lowest average tracking error clearly demonstrate that our approach surpasses other state-of-the-art methods. We present our RGB-D tracking performance on the real-world ScanNet dataset [40] in Tab. II. Compared to the results on the Replica dataset, all methods exhibit degraded performance due to the noisy and sparse depth sensor input, as well as lower color image quality caused by motion blur. Specifically for semantic understanding, the semantic annotations provided by ScanNet are significantly noisier than those in Replica, containing noticeable noise and imprecise boundaries. All of these issues negatively impact the overall performance of semantic SLAM systems. We evaluate all six sequences and show that our method achieves performance comparable to state-of-the-art approaches [25, 16], demonstrating the robustness of our proposed method under these challenging conditions. The tracking performance on the TUM-RGBD dataset is shown in Tab. III. Since the TUM-RGBD dataset does not contain semantic ground truth, we evaluate our method based on tracking accuracy here. Despite the low-quality RGB and depth inputs, as well as the absence of semantic information, all of which could negatively impact overall performance, our method still achieves similar performance to the state-of-the-art approach [16], with a localization error of less than $0.2$ cm, while operating at a significantly faster runtime. A detailed analysis of the runtime performance will be provided in the following section.
@@
IV-C Global Mapping Performance
Following previous methods [25, 23, 24], we evaluate the depth error on the Replica dataset [41] to assess overall mapping performance, as it provides the highest-quality depth maps for comparison. With an accurate understanding of semantics, our method effectively learns precise global geometry in unknown environments, benefiting both tracking and mapping in the semantic SLAM system. Our method achieves superior results in 7 out of 8 sequences and maintains a low average depth error across all sequences, with a difference of approximately $0.1$ cm compared to the best performing. These results demonstrate that our approach have strong semantic mapping capability, highlighting our superior mapping performance.
IV-D Rendering Quality
We present the quantitative analysis of the rendering performance of our proposed method in Tab. V. Following previous works, Point-SLAM [45] and NICE-SLAM [33], the evaluation is conducted on the input views. As shown in Tab. V, our method achieves superior image rendering performance compared to state-of-the-art approaches across all 8 sequences, as well as in the overall average evaluation. As discussed in previous sections, incorporating semantic information introduces a significant computational burden to the joint optimization process within the SLAM system. However, our method demonstrates notably improved rendering performance compared to all existing methods, including the non-semantic SLAM approaches. Although the additional computation for semantic learning increases the optimization complexity, the accurate semantic understanding enhances both geometric and appearance learning through joint optimization, effectively improving rendering performance.
The visual comparison of rendered images is presented in Fig. 6. Benefiting from precise semantic information learning and the improved SLAM system, the rendering results exhibit significantly improved detail preservation, such as more accurate and well-defined lamp boundaries in the first and fifth columns, and precise details even in fine structures like shutters in the second column. Moreover, our method produces high-quality renderings with fewer noise artifacts, as evidenced by the improved rendering quality of the floor in the fourth column and the chair in the sixth column. This demonstrates that better mapping performance leads to enhanced rendering quality. The joint optimization of semantic losses and photometric loss is mutually beneficial, facilitating more accurate global 3D reconstruction through semantic learning, which in turn improves the rendering performance of our Hier-SLAM++.
TABLE IV: Reconstruction performance Depth L1 (cm) on Replica. Best results are highlighted as first, second.
@@ IV-D Rendering Quality
IV-E Running time
The runtime results for the compared methods are provided in Tab. VI, covering both NeRF-based and Gaussian Splatting-based SLAM methods. Since the compared baseline using a flat representation (Hier-SLAM++-flat) cannot run successfully on an NVIDIA RTX 4090 due to GPU memory limitations, we report the runtime separately for RTX 4090 and L40S. In the first block of Tab. VI, we observe that, compared to NeRF-based methods [33], Gaussian-based state-of-the-art methods [16] achieve significantly higher operating speeds, benefiting from the fast rendering and efficient optimization capabilities inherent to the new 3D representation, i.e., Gaussian Splatting. Furthermore, with the carefully optimized forward and backward Gaussian rendering module introduced in our system, which serves as the core component of Gaussian Splatting-based SLAM systems, our method (Hier-SLAM++- w/o sem) achieves up to 2.4× faster tracking and 2.2× faster mapping compared to the best-performing method [16].
In the second block of Tab. VI, when integrating semantic information, our method maintains high efficiency by leveraging hierarchical semantic coding (Hier-SLAM++-one-hot), achieving nearly 3× faster tracking compared to flat semantic coding (Hier-SLAM++-flat). Compared to the state-of-the-art method [25], our Hier-SLAM++ attains similar operational speed but significantly outperforms in semantic segmentation accuracy. Thanks to the improved semantic decoder in our hierarchical semantic loss, our method achieves a $13\%$ mIoU improvement while incurring only an additional $0.04$ s per frame for tracking and $2$ s per frame for mapping. Further details on this improvement will be provided in the next section. Additionally, our binary version (Hier-SLAM++-binary) is able to achieve more efficient operation performance. Notably, our Hier-SLAM++ achieves a rendering speed of 2000 FPS, which increases to 3000 FPS when semantic information is not included.
TABLE V: Rendering performance PSNR, SSIM, LPIPS on Replica. Best results are highlighted as first, second.

### v2

@@ III-C Monocular setting with geometric priority
Accurate depth information provides valuable geometric cues for global map reconstruction and pose estimation in SLAM systems. While RGB-D methods directly acquire depth from sensors, monocular SLAM approaches [28, 29, 30, 31] must rely on multi-view triangulation to infer depth and initialize mapping, which increases the difficulty of simultaneously performing localization and mapping. To address this limitation, a number of methods [71, 72, 73, 74, 75, 76] introduce depth priors from learned depth networks into SLAM systems, achieving promising results. More recently, a new trend has emerged: 3D feed-forward models, which can generate accurate global 3D representations of previously unseen environments after training. In our work, we adopt a precise and highly generalizable 3D feed-forward method to obtain the geometric prior used in monocular SLAM. Specifically, we employ DUSt3R [24], which produces both 3D point clouds and pose estimations from sparse-view image inputs, serving as a strong geometric prior in our system. Details of the monocular SLAM setting in Hier-SLAM++ are provided in the following section.
@@ III-C1 Geometric prior acquisition
Each image set is then fed into DUSt3R to generate geometric information, specifically the depth maps. While DUSt3R is also capable of estimating camera poses for each input frame, our experiments reveal that its pose predictions exhibit relatively low accuracy, often performing worse than a simple constant-velocity pose model when integrated into a SLAM pipeline. Therefore, we retain only the geometric (depth) information predicted by DUSt3R, as it provides more reliable guidance for our SLAM system.
@@ III-C2 Geometric prior correction
The depth predicted by DUSt3R is scale-ambiguous and lacks metric consistency across image sets. This leads to differences in scale between the predicted geometry and the SLAM system’s estimated global map, which evolves over time. We therefore incorporate a geometric prior correction step to resolve this scale inconsistency.
To align the scale of the depth prior with the current global map, we first estimate the scale and shift parameters by comparing the rendered depth from the SLAM map with the predicted depth from DUSt3R. Specifically, for each input frame $I_{c}$, we render a depth map $D_{c}$ and silhouette map $S_{c}$ from the global 3D Gaussian map $G$ using the current camera pose $\mathbf{T}_{c}$:
@@
where $\mathcal{R}$ represents the rasterization procedure for the 3D Gaussian map. We then compute the global scale and shift parameters $\lambda_{c}$ and $\tau_{c}$ by aligning the DUSt3R-predicted depth $\hat{D}c$ to the rendered depth $D_{c}$ using least squares estimation over the visible regions $M_{c}$, as defined by the silhouette map $S_{c}$:
@@
where $\hat{D}_{c}^{\text{align}}=\lambda_{c}\cdot\hat{D}_{c}+\tau_{c}$ denotes the aligned depth. This simple affine transformation is sufficient to resolve the scale ambiguity inherent in the geometric priors obtained from DUSt3R, while preserving their relative 3D structure for integration into the monocular SLAM pipeline.
Secondly, the calculated scale and shift parameters are used as initial values and are jointly optimized with the 3D Gaussians during the map optimization process, guided by a depth supervision loss:
$\{\lambda_{c},\tau_{c},G\}=\arg\min\big{(}|\hat{D}_{c}^{\text{align}}-D_{c}|M_{c}+L_{\text{reg}}\big{)}$ (17)
where $L_{\text{reg}}$ is a regularization term defined as the absolute deviation between the optimized parameters and their initialization, preventing overfitting or convergence to poor local minima. By incorporating this optimization into the SLAM operation, the depth prior can be continuously refined online as new observations arrive, effectively providing geometric supervision throughout the SLAM process. This not only improves the quality of the global map reconstruction but also enhances the accuracy of camera pose estimation.
@@ III-D Semantic Gaussian Splatting SLAM
The global semantic 3D Gaussian Splatting SLAM system is illustrated in Fig. 2. Built upon the semantic 3D Gaussian representation, global reconstruction (Mapping) and camera pose estimation (Tracking) are alternately performed over the input video stream. In the RGB-D setting, depth information for each frame is directly acquired from the depth sensor. In the monocular setting, a feed-forward model is employed to generate geometric priors, as described in the previous section. The construction of the semantic 3D Gaussians and the overall semantic SLAM system are detailed in the following sections.
@@ III-D1 Semantic 3D Gaussian representation
Our semantic 3D Gaussian representation encodes both geometric and semantic information using Gaussian primitives augmented with hierarchical semantic embeddings. Motivated by [16], we adopt isotropic, view-independent Gaussians and integrate them with semantic feature vectors. Each semantic Gaussian is parameterized by its color $\bm{c}$, center position $\bm{\mu}$, radius $r$, opacity $o$, and semantic embedding $\bm{h}$. The influence of each Gaussian $G(\bm{X})$ [13] at 3D point $\bm{X}$ follows the standard Gaussian formulation:
@@
Following [13], every semantic 3D Gaussian primitive within the current global map is projected onto the image plane using the tile-based differentiable $\alpha$-compositing rendering. The semantic map $H$ is rendered as follows:
$H=\sum_{i=1}^{n}\bm{h}_{i}\alpha_{i}(\bm{X})T_{i}\quad\text{with}\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{X}))$ (19)
@@
Different from previous work [16], which uses separate forward and backward Gaussian modules for different parameters, our Hier-SLAM++ system adopts a unified forward and backward modules that processes all parameters, including semantics, color, depth, and silhouette images, which significantly improving overall runtime efficiency.
@@ III-D2 Tracking
$\mathbf{T}_{c}=\arg\min_{\mathbf{T}_{c}}L_{\text{Track}}(G,\mathbf{T}_{c})$ (21)
@@ III-D3 Mapping
The global map information, including semantic information, is optimized in the mapping procedure while keeping camera poses $\mathbf{T}_{i}$ fixed:
$G=\arg\min_{G}L_{\text{Map}}(G,\mathbf{T}_{i})$ (23)
@@ IV-A Experiment settings
The experiments are conducted on both synthetic and real-world datasets, including ScanNet [77], Replica [78], and TUM-RGBD [79]. Following the evaluation metrics which used in previous SLAM methods [16, 59], we leverage ATE RMSE (cm) to quantify SLAM tracking accuracy. For mapping evaluation, Depth L1 (cm) is used to measure accuracy. To assess image rendering quality, we adopt PSNR (dB), SSIM, and LPIPS as evaluation metrics. In line with previous works [68, 7, 11], we assess semantic rendering performance using mIoU (mean Intersection over Union across all classes) to represent global semantic information. To evaluate computational efficiency, we report the runtime of the proposed SLAM method. For comparison, we benchmark our approach against state-of-the-art dense visual SLAM methods, including both NeRF-based and 3D Gaussian SLAM approaches, to demonstrate its effectiveness. Additionally, we compare with state-of-the-art semantic SLAM methods, covering both NeRF-based and Gaussian-based techniques, to highlight its capability in hierarchical semantic representation and scalability. Experiments are conducted on an NVIDIA L40S GPU, while runtime tests are evaluated on both the L40S and RTX 4090 to enable broader comparisons.
For experimental settings, the semantic embedding of each Gaussian primitive is initialized randomly. The semantic optimization loss weights $\omega_{1}$ and $\omega_{2}$ are set to $1.0$ and $0.0$, respectively, for the first $\eta$ iterations, where $\eta$ is set as $15$. Afterwards, $\omega_{1}$ remains $1.0$, while $\omega_{2}$ is increased to $5.0$. This strategy means that we first apply the Inter-level loss to initialize the hierarchical semantic coding, followed by incorporating the Cross-level loss to refine the whole embedding. For tracking loss, we use $\delta=0.99$, with weights $w_{1}=1.0$ and $w_{2}=0.5$. For mapping, the loss weights are set as follows: $w_{3}=1.0$, $w_{4}=0.5$, and $w_{5}=0.2$. The two types of semantic embeddings proposed in our method, one-hot and binary, are evaluated and reported across multiple experiments, including tracking, mapping, rendering, runtime, and semantic understanding, providing a comprehensive ablation study on the effectiveness and compactness of the hierarchical semantic representation.
@@
Best results are highlighted as first, second, third.
@@ IV-B Camera Tracking Accuracy
Tab. I presents our tracking performance on the Replica dataset [78] in the RGB-D setting, compared with state-of-the-art dense SLAM methods, both with and without semantic information. By accurately learning semantic information, our Hier-SLAM++ outperforms all existing methods in 6 out of 8 sequences. For the remaining two sequences, our method achieves nearly the same performance with sub-millimeter level localization error. The lowest average tracking error clearly demonstrate that our approach surpasses other state-of-the-art methods. We present our RGB-D tracking performance on the real-world ScanNet dataset [77] in Tab. II. Compared to the results on the Replica dataset, all methods exhibit degraded performance due to the noisy and sparse depth sensor input, as well as lower color image quality caused by motion blur. Specifically for semantic understanding, the semantic annotations provided by ScanNet are significantly noisier than those in Replica, containing noticeable noise and imprecise boundaries. All of these issues negatively impact the overall performance of semantic SLAM systems. We evaluate all six sequences and show that our method achieves performance comparable to state-of-the-art approaches [25, 16], demonstrating the robustness of our proposed method under these challenging conditions. The tracking performance on the TUM-RGBD dataset is shown in Tab. III. As this dataset does not provide semantic ground truth, we evaluate our method solely based on tracking accuracy. Without access to semantic information, our system operates as a standard visual SLAM pipeline. As a result, the tracking accuracy becomes comparable to the state-of-the-art method [16] in the TUM-RGBD dataset.
@@
TABLE IV: Reconstruction performance Depth L1 (cm) on Replica. Best results are highlighted as first, second, third.
@@
IV-C Global Mapping Performance
Following prior work [25, 67, 68], we evaluate depth accuracy on the Replica dataset [78], which provides high-quality ground truth depth maps for reliable evaluation of overall mapping performance. The quantitative results are presented in Tab. IV. Our method achieves the lowest depth error in 7 out of 8 sequences and maintains a consistently low average error across all sequences, with only a $0.1$ cm difference compared to the best-performing baseline. By leveraging semantic understanding, our approach enables the learning of precise geometry in unseen environments, which in turn benefits both tracking and mapping within the whole semantic SLAM system. These results confirm the strong semantic mapping capability of our method and highlight its superior performance in dense 3D reconstruction.
IV-D Rendering Quality
Rendering performance is becoming an increasingly important evaluation metric, particularly for SLAM systems built upon new scene representations, such as neural implicit representations and gaussian splatting. Following prior works [62, 59], the evaluation is conducted on the input views of the Replica dataset. We present a quantitative analysis of the rendering performance of our proposed method in Tab. V. Our method outperforms state-of-the-art approaches in rendering quality across all 8 sequences, as well as in the overall average, surpassing both semantic and non-semantic SLAM baselines. The accurate semantic understanding learned by our system enhances both geometric and appearance representations through joint optimization, further improving the overall rendering fidelity of the SLAM system.
The visual comparison of rendered images is shown in Fig. 4. Benefiting from accurate semantic understanding and an improved SLAM framework, our method achieves significantly better detail preservation. Notable improvements include sharper and more accurate lamp boundaries in the first and fifth columns, as well as fine structural details, such as the shutters in the second column. Furthermore, our method produces high-quality renderings with fewer noise artifacts, evident in the clearer floor in the fourth column and the cleaner reconstruction of the chair in the sixth column. These results demonstrate that improved mapping quality directly contributes to enhanced rendering fidelity. The joint optimization of semantic and photometric losses proves mutually beneficial: semantic learning facilitates more accurate global 3D reconstruction, which in turn leads to better rendering performance in our Hier-SLAM++ system.
Figure 4: Visualization of RGB image rendering performances on the Replica dataset [78]. The first to third rows show the rendering outputs of the comparison methods. The fourth row presents the image rendering results of our Hier-SLAM++. The last row displays the ground truth images for reference. We use red boxes to highlight key differences regions within the image. The visual comparison clearly demonstrates that our method achieves superior rendering quality, closely resembling the ground truth images compared to state-of-the-art approaches.
TABLE V: Rendering performance PSNR, SSIM, LPIPS on Replica. Best results are highlighted as first, second, third.